# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 3.2. Affinization of open Kulikov surfaces
In this section, we construct for every open Kulikov surface $\mathcal{X}_{0}$ a proper surjective morphism $f_{0}\colon\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ to a non-normal affine surface $\overline{\mathcal{X}}_{0}$. We show that $f_{0}$ is the affinization of $\mathcal{X}_{0}$, and that the affine surface $\overline{\mathcal{X}}_{0}$ is a degenerate cusp singularity.
Let
$\mathcal{X}_{0}=\bigcup_{v\in P_{\mathbb{Z}}}X^{v}$
be an open Kulikov surface, with tropicalization $(P,\mathscr{P})$, where we denote by $X^{v}$ the irreducible component of $\mathcal{X}_{0}$ corresponding to the integral points $v\in P_{\mathbb{Z}}$.
We denote by $P_{\mathbb{Z}}(0)$ (resp. $P_{\mathbb{Z}}(1)$ and $P_{\mathbb{Z}}(2)$) the set of integral points $v\in P_{\mathbb{Z}}$ such that $X^{v}$ is a 0-surface (resp. a 1-surface and a 2-surface). By Definition 3.7 ii), for every boundary integral point $v\in(\partial P)_{\mathbb{Z}}$, we have either $v\in P_{\mathbb{Z}}(0)$ or $v\in P_{\mathbb{Z}}(1)$. By Definition 3.7 i), $P$ is topologically a disk, and so the boundary $\partial P$ is topologically a circle. If one fixes an orientation of $\partial P$, we have a corresponding cyclic orientation of the boundary vertices $v\in(\partial P)_{\mathbb{Z}}=P_{\mathbb{Z}}(0)\cup P_{\mathbb{Z}}(1)$. We denote by $(v_{i})_{i\in\mathbb{Z}/N\mathbb{Z}}$ the cyclically ordered vertices in $P_{\mathbb{Z}}(0)$. Moreover, $P_{\mathbb{Z}}(2)$ is exactly the set of vertices of $\mathscr{P}$ contained in the interior of $P$.
@@
If $v\in P_{\mathbb{Z}}(2)$, then $X^{v}$ is compact, and so $(\overline{X}^{v},\partial\overline{X}^{v})$ is just a point. If $v\in P_{\mathbb{Z}}(1)$, then $(\overline{X}^{v},\partial\overline{X}^{v})=(\mathbb{A}^{1},\mathbb{A}^{1})$ by Definition 3.4. Finally, for $i\in\mathbb{Z}/N\mathbb{Z}$, corresponding to an integral point $v_{i}\in P_{\mathbb{Z}}(0)$, the affinization $(\overline{X}^{v_{i}},\partial\overline{X}^{v_{i}})$ is an affine toric surface by Definition 3.6. Moreover, $\partial\overline{X}^{v_{i}}$ is the corresponding toric divisor, that is, the union of two irreducible toric divisors $(\partial\overline{X}^{v_{i}})_{i-1}$ and $(\partial\overline{X}^{v_{i}})_{i+1}$ corresponding to the two boundary edges of $\mathscr{P}$ adjacent to $v_{i}$ and contained in the intervals $[v_{i-1},v_{i}]$ and $[v_{i},v_{i+1}]$ of $\partial P$ respectively. Both $(\partial\overline{X}^{v_{i}})_{i-1}$ and $(\partial\overline{X}^{v_{i}})_{i+1}$ are isomorphic to $\mathbb{A}^{1}$, with 00 corresponding to the torus fixed point of $\overline{X}^{v_{i}}$.
We define
$\overline{\mathcal{X}}_{0}:=\bigcup_{v\in P_{\mathbb{Z}}(0)}\overline{X}^{v}=\bigcup_{i\in\mathbb{Z}/N\mathbb{Z}}\overline{X}^{v_{i}}\,,$ (3.1)
where for every $i\in\mathbb{Z}/N\mathbb{Z}$, we identify $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$ along their toric divisors $(\partial\overline{X}^{v_{i}})_{i+1}\simeq\mathbb{A}^{1}$ and $(\partial\overline{X}^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$ by the isomorphism induced by the gluing isomorphism between the irreducible components $(\partial{X}^{v_{i}})_{i+1}$ and $(\partial{X}^{v_{i+1}})_{i}$ of the double locus $\mathcal{X}_{0}$ corresponding to the two boundary edges of $\mathscr{P}$ adjacent to $v_{i}$. In particular, this isomorphism preserves $0\in\mathbb{A}^{1}$, and so identifies the torus fixed points of $\overline{X}^{v_{i}}$ and $\overline{X}^{v_{i+1}}$.
@@
By construction, the irreducible components of $\overline{\mathcal{X}}_{0}$ are the the affine toric surfaces $\overline{X}^{v_{i}}$, and so $\overline{\mathcal{X}}_{0}$ is itself affine by [51, Exercise III.3.2]. Moreover, the double locus of $\overline{\mathcal{X}}_{0}$ is a union of copies of affine lines $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$, and all intersecting at a single common point $x_{0}\in\overline{\mathcal{X}}_{0}$, referred below as the central point of $\overline{\mathcal{X}}_{0}$, see Figure 3.1. It follows that $\overline{\mathcal{X}}_{0}$ is a degenerate cusp singularity as in [86, §1], and so is in particular Gorenstein semi-log-canonical by [68, Theorem 4.21].
[…]
Figure 3.1. Affinization $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ of an open Kulikov surface $\mathcal{X}_{0}$.
Finally, we construct a morphism $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ as follows. We first define a morphism $f^{v}:X^{v}\rightarrow\overline{\mathcal{X}}_{0}$ for every integral point $v\in P_{\mathbb{Z}}$. To do that, we distinguish the three cases of 00-, $1$-, and $2$-surface:
  • (i)
    If $v\in P_{\mathbb{Z}}(0)$, we define $f^{v}$ as the composition of the affinization map $X^{v}\rightarrow\overline{X}^{v}$ of the 00-surface $X^{v}$ with the inclusion $\overline{X}^{v}\subset\overline{\mathcal{X}}_{0}$ of $\overline{X}^{v}$ as an irreducible component of $\overline{\mathcal{X}}_{0}$.
  • (ii)
    If $v\in P_{\mathbb{Z}}(1)$, contained in the interval $[v_{i},v_{i+1}]$ in $\partial P$, we define $f^{v}$ as the composition of the affinization map $X^{v}\rightarrow\overline{X}^{v}\simeq\mathbb{A}^{1}$ of the $1$-surface $X^{v}$ with the inclusion $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}\subset\overline{\mathcal{X}}_{0}$ of the irreducible component $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$ of the double locus of $\overline{\mathcal{X}}_{0}$.
  • (iii)
    If $v\in P_{\mathbb{Z}}(2)$, that is, $v\notin(\partial P)_{\mathbb{Z}}$, we define $f^{v}$ as the constant map contracting the compact $2$-surface $X^{v}$ to the central point $x_{0}\in\overline{\mathcal{X}}_{\mathscr{P},0}$.
By construction, the morphisms $f^{v}$ and $f^{v^{\prime}}$ agree on the intersections $X^{v}\cap X^{v^{\prime}}$ in $\mathcal{X}_{0}$, and so they glue together to produce a morphism $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$.
@@
The morphism $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ is a contraction, that is, a surjective proper morphism with $(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}=\mathcal{O}_{\overline{\mathcal{X}}_{0}}$. Moreover, we have $R^{i}(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}=0$ for all $i>0$.

### v3

@@ 3.2. Affinization of open Kulikov surfaces
Let $\mathcal{X}_{0}=\bigcup_{v\in P_{\mathbb{Z}}}X^{v}$ be an open Kulikov surface, with tropicalization $(P,\mathscr{P})$, where we denote by $X^{v}$ the irreducible component of $\mathcal{X}_{0}$ corresponding to the integral point $v\in P_{\mathbb{Z}}$. Denote by $\mathcal{X}_{0}^{\mathrm{can}}$ the affinization of $\mathcal{X}_{0}$, that is, $\mathcal{X}_{0}^{\mathrm{can}}=\mathrm{Spec}\,H^{0}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})$, and by $f_{0}:\mathcal{X}_{0}\rightarrow\mathcal{X}_{0}^{\mathrm{can}}$ the affinization morphism.
Figure 5. Affinization $f_{0}:\mathcal{X}_{0}\rightarrow\mathcal{X}^{\mathrm{can}}_{0}$ of an open Kulikov surface $\mathcal{X}_{0}$.
Lemma 3.13.
Let $\mathcal{X}_{0}$ be an open Kulikov surface. Then, the affinization morphism $f_{0}:\mathcal{X}_{0}\rightarrow\mathcal{X}^{\mathrm{can}}_{0}$ is a contraction, that is, a surjective proper morphism with $(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}=\mathcal{O}_{\mathcal{X}^{\mathrm{can}}_{0}}$. Moreover, we have $R^{i}(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}=0$ for all $i>0$.
Proof.
We denote by $P_{\mathbb{Z}}(0)$ (resp. $P_{\mathbb{Z}}(1)$ and $P_{\mathbb{Z}}(2)$) the set of integral points $v\in P_{\mathbb{Z}}$ such that $X^{v}$ is a 0-surface (resp. a 1-surface and a 2-surface). By Definition 3.7 ii), we have $P_{\mathbb{Z}}(0)\neq\emptyset$, and, for every boundary integral point $v\in(\partial P)_{\mathbb{Z}}$, we have either $v\in P_{\mathbb{Z}}(0)$ or $v\in P_{\mathbb{Z}}(1)$. By Definition 3.7 i), $P$ is topologically a disk, and so the boundary $\partial P$ is topologically a circle. If one fixes an orientation of $\partial P$, we have a corresponding cyclic orientation of the boundary vertices $v\in(\partial P)_{\mathbb{Z}}=P_{\mathbb{Z}}(0)\cup P_{\mathbb{Z}}(1)$. We denote by $(v_{i})_{i\in\mathbb{Z}/N\mathbb{Z}}$ the cyclically ordered vertices in $P_{\mathbb{Z}}(0)$. Moreover, $P_{\mathbb{Z}}(2)$ is exactly the set of vertices of $\mathscr{P}$ contained in the interior of $P$.
@@
If $v\in P_{\mathbb{Z}}(2)$, then $X^{v}$ is compact, and so $(\overline{X}^{v},\partial\overline{X}^{v})$ is just a point. If $v\in P_{\mathbb{Z}}(1)$, then $(\overline{X}^{v},\partial\overline{X}^{v})=(\mathbb{A}^{1},\mathbb{A}^{1})$ by Definition 3.4. For $i\in\mathbb{Z}/N\mathbb{Z}$, corresponding to an integral point $v_{i}\in P_{\mathbb{Z}}(0)$, the affinization $(\overline{X}^{v_{i}},\partial\overline{X}^{v_{i}})$ is an affine toric surface by Definition 3.6. Moreover, $\partial\overline{X}^{v_{i}}$ is the corresponding toric divisor, that is, the union of two irreducible toric divisors $(\partial\overline{X}^{v_{i}})_{i-1}$ and $(\partial\overline{X}^{v_{i}})_{i+1}$ corresponding to the two boundary edges of $\mathscr{P}$ adjacent to $v_{i}$ and contained in the intervals $v_{i-1},v_{i}]$ and $v_{i},v_{i+1}]$ of $\partial P$ respectively. Both $(\partial\overline{X}^{v_{i}})_{i-1}$ and $(\partial\overline{X}^{v_{i}})_{i+1}$ are isomorphic to $\mathbb{A}^{1}$, with 0 corresponding to the torus fixed point of $\overline{X}^{v_{i}}$.
@@
The irreducible components of $\mathcal{X}^{\mathrm{can}}_{0}$ are the affine toric surfaces $\overline{X}^{v_{i}}$ for $i\in\mathbb{Z}/N\mathbb{Z}$. Moreover, the double locus of $\mathcal{X}^{\mathrm{can}}_{0}$ is a union of copies of affine lines $(\partial X^{v_{i}})_{i+1}\simeq(\partial X^{v_{i+1}})_{i}\simeq\mathbb{A}^{1}$, and all intersecting at a single common point $x_{0}\in\mathcal{X}^{\mathrm{can}}_{0}$, referred below as the central point of $\mathcal{X}^{\mathrm{can}}_{0}$, see Figure 5.
$0\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}\mathcal{O}_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}\mathcal{O}_{X^{f}}\rightarrow 0\,$, (3.1)
as in the proof of Lemma 3.10. We have the following vanishing results:
  • i)
  • ii)
  • iii)
  • iv)
Hence, the first page of the spectral sequence obtained by applying $R(f_{0})_{\star}$ to the complex (3.1) reduces to the complex
$0\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}(f_{0})_{\star}\mathcal{O}_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}(f_{0})_{\star}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}(f_{0})_{\star}\mathcal{O}_{X^{f}}\rightarrow 0\,.$
The stalk of this complex at the point $x_{0}$ is isomorphic to the tensor product of the local ring $\mathcal{O}_{\mathcal{X}_{0}^{\mathrm{can}},x_{0}}$ with the simplicial chain complex defined by the polyhedral decomposition $\mathscr{P}$ of the tropicalization $P$. Since $P$ is topologically a disk, we have $H^{0}(P,\mathbb{Z})=\mathbb{Z}$ and $H^{i}(P,\mathbb{Z})=0$ for $i>0$, and so the stalk of $R^{i}(f_{0})_{\star}\mathcal{O}_{\mathcal{X}_{0}}=0$ at $x_{0}$ is equal to zero for $i>0$. Similarly, at a point of the double locus (resp. a smooth point) of $\mathcal{X}_{0}^{\mathrm{can}}$, the disk is replaced by an interval (resp. a point), and so Lemma 3.13 follows. ∎
@@
Let $\mathcal{X}_{0}$ be an open Kuliukov surface. Then, the affine surface $\mathcal{X}_{0}^{\mathrm{can}}$ is a degenerate cusp singularity as in [90, §1]. In particular, $\mathcal{X}_{0}^{\mathrm{can}}$ is a Gorenstein semi-log-canonical surface with $\omega_{\mathcal{X}_{0}^{\mathrm{can}}}=\mathcal{O}_{\mathcal{X}_{0}^{\mathrm{can}}}$.